OpenStudy (anonymous):

Two stores sell the same television for the same original price. Store A advertises that the television is on sale for 30% off the original price. Store B advertises that it is reducing the television���������s price by $250. When Allison compares the sale prices of the television in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the television���������s original price.

OpenStudy (texaschic101):

.70p = p - 250 -- 30% off of origjnal price means you are actually paying 70% of original price (p)....which equals original price (p) - 250 subtract p from both sides .70p - p = -250 -.30p = -250 -- divide both sides by -0.30 p = -250/-0.30 p = 833.33 check.. store A....70% of 833.33 = .70(833.33) = 583.33 store B...p - 250 = 833.33 - 250 = 583.33 so original price is $ 833.33

OpenStudy (anonymous):

A. 0.3p = p - 250 B. 0.7p = p - 250 C. 0.3p = p + 250 D. 0.7(p - 250) = p